Fin two consectuive odd integers such that 4 times smaller the number is 7 more than 3 times the larger number.
Answer by checkley77(12569)

Let x & x+2 be the 2 integers.
4x=3(x+2)+7
4x=3x+6+7
4x-3x=13
x=13 ans. for the smaller integer.
13+2=15 ans. for the larger integer.
Proof:
4*13=3*15+7
52=45+7
52=52